Comparing Three Depreciation Methods Dexter Industries purchased packaging equipment on January 8 for $245,200. The equipment was expected to have a useful life of three years, or 7,500 operating hours, and a residual value of $20,200. The equipment was used for 3,000 hours during Year 1, 2,325 hours in Year 2, and 2,175 hours in Year 3. Required: 1. Determine the amount of depreciation expense for the three years ending December 31, Year 1, Year 2, Year 3, by (a) the straight-line method, (b) the units-of-activity method, and (c) the double-declining-balance method. Also determine the total depreciation expense for the three years by each method. Note: For all methods, round the answer for each year to the nearest whole dollar. Depreciation Expense Year Straight-Line Method Units-of-Activity Method Double-Declining-Balance Method Year 1 $ $ $ Year 2 $ $ $ Year 3 $ $ $ Total $ $ $ 2. What method yields the highest depreciation expense for Year 1? 3. What method yields the most depreciation over the three-year life of the equipment?
Answer #1

1) Calculate depreciation expense:

a) Straight line method :

Depreciation expense = (245200-20200/3) = $75000 per year

Depreciation expense
Year 1 75000
Year 2 75000
Year 3 75000
Total 225000

b) Unit of activity method :

Depreciation rate = (245200-20200/7500) = $30 per hour

Depreciation expense
Year 1 3000*30 = 90000
Year 2 2325*30 = 69750
Year 3 2175*30 = 65250
Total 225000

c) Double decline balance :

Dep rate = 100/3*2 = 66.67%

Depreciation expense
Year 1 245200*2/3 = 163467
Year 2 245200*1/3*2/3 = 54489
Year 3 7044
Total 225000

2) Highest depreciation expense for year 1 = Double decline balance method

3) All three methods